40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Feb 17 23:50:55 2011, tara, DailyProgress, VCO, PDH box irregular gain behavior PDH_input_refernoise.pngACAVTF.pngnoise.png
    Reply  Fri Feb 18 23:52:42 2011, tara, DailyProgress, VCO, PDH box irregular gain behavior VCOTF.pngrefernoise.pngbeat_2011_02_18.png
Message ID: 503     Entry time: Thu Feb 17 23:50:55 2011     Reply to this: 504
Author: tara 
Type: DailyProgress 
Category: VCO 
Subject: PDH box irregular gain behavior 

I measured the input referred noise of the VCO's PDH box at different gain setting.

The input refer noise varies with gain setup. The result is plotted below


When both cavities are in lock, the VCO receives the feedback signal from the PDH box to control the laser frequency.

The PSD of the feedback signal to VCO varies with gain setup on the PDH box significantly, so it needs to be checked.

To decide the gain setup range for test, the gain is adjusted while the beat noise is observed. The minimum gain and maximum gain

where the beat noise still looks nice are chosen, which are 4 and 7 as read on the knob dial.


Then the noise of the PDH box is measured at different gain setups (4,5,6,7.) When

1) the PDH input is shorted (for PDH intrinsic noise alone ), and

2) the beam on the RFPD is blocked (for whole set of RFPD, mixer and PDH box)

 (data from gain 4 with RFPD has to be remeasured )


Then the open loop TF is measured. The measured noise divided by the TF is the input referred noise of our system.

The TF in the measured region can be approximated with 1/f slope.

 The input referred goes down alot from gain 4 to 5, and not much from 5 to 7;


The input referred noise doesn't seem correct. I'll double check it later.



Attachment 1: PDH_input_refernoise.png  46 kB  | Hide | Hide all
Attachment 2: ACAVTF.png  32 kB  | Hide | Hide all
Attachment 3: noise.png  49 kB  | Hide | Hide all
ELOG V3.1.3-